a) Find the area of a trapezoid with a height of 6 m and bases of 14 m and
21 m
b) Find the perimeter of the parallelogram below.

9inches for the top
5 inches for the side
4 inches ----cut down the middle of trapezoid



Answer by rfer(16322)

1)
A=1/2h(a+b)
A=1/2(6)(14+21)
A=3(35)
A=105sqm
2)
P=2b+2s
P=2*9+2*5
P=18+10
P=28in